OpenStudy (anonymous):

Find x and y such that the points A(-1,3), B(x,9) and C(5,y) are collinear with point B lying midway between Points A and C.

OpenStudy (ghazi):

use mid point formula that states that mid point is mean of two end points \[\frac{ -1+5 }{ 2 }=x\] similarly for y

OpenStudy (ghazi):

\[\frac{ 3+y }{ 2 }=9\]